# Rethinking the Vulnerabilities of Face Recognition Systems: From a Practical Perspective

### v1

@@ 3.3 Experimental Exploration and Analysis
$\mathrm{cos}(f_{\theta}(\Phi(x_{q})),e_{q})\leq\delta_{match}\quad or$ (3)
$\mathrm{cos}(f_{\theta}(\Phi(x_{q})),e_{k})\geq\delta_{match},k\neq q$
where faical features $e_{q},e_{k}\in\mathbb{D}$ and $\Phi(\cdot)$ denotes the transformation that makes $f_{\theta}(\Phi(x_{q}))$ different from $e_{q}$, leading to the misidentification. In practice, $\Phi(\cdot)$ simulates the factors introduced by the environment (e.g., light intensity, quality of the images, etc) and human (makeup, shooting angles, expressions, decorations, etc). All of these factors work together and result in recognition errors.
@@
where $l(\cdot)$ is the loss function for optimization. We also restrict the changing area of $p$ with $\|p\|_{n}\leq\gamma$. The details of the algorithm can be found in Appendix A. For the sake of experimental simplicity, we only set the changing areas of $p$ to 10% (for image space), and $p$ is randomly initialized in each iteration to avoid its influence on the optimization of $m$. The result is shown in Figure 4, using CelebA [35] and MobileFace [36] for optimization. From the similarity between $x$ and $x\oplus p$, we can indicate that some features are more essential for $f_{\theta}$ than others, and only a small part of the occlusion of these key features can blind FRS. Based on the analysis above, the unrecognition ($\mathrm{cos}(f_{\theta}(\Phi(x_{q})),e_{q})\leq\delta_{match}$) of FRS can be attributed to the influence of various factors on the key features of feature extraction.
@@
The above inference can also be associated with the misidentification ($\mathrm{cos}(f_{\theta}(\Phi(x_{q})),e_{k})\geq\delta_{match},k\neq q$) of FRS, since faces without distinguished features could easily be matched with an ”average face” which also lacks distinguished features in the database. To validate this conjecture, we randomly sample 1000 face pairs $(x_{i},x_{j})$ from CelebA as $\mathbb{D}_{pair}$, where the identity of $x_{i}$ is different with $x_{j}$. Then, we use prefabricated masks (see in Appendix A) $\mathbb{M}=\{m_{0},m_{1},...,m_{i}\}$ for different parts of the face. Based on the above preparations, we can draw a histogram of the similarity of $\mathrm{cos}(f_{\theta}(x_{i}\cdot(1-m_{k})),f_{\theta}(x_{j}\cdot(1-m_{k})))$, for $m_{k}\in\mathbb{M}$ and $(x_{i},x_{j})\in\mathbb{D}_{pair}$. Figure 5 illustrates that the distribution of face similarity will move forward as a result of the addition of the masks, and different regions present different impacts on the similarity changes. For instance, mask regions, including the eye and nose, can largely increase the similarity of faces from different identities. However, masks of mouth or cheek have less impact on the feature extraction of FRS. With these observations, we can also arrive at a conclusion consistent with the previous experiments, which is that facial feature extractors tend to focus more on the local features of faces to distinguish between facial images from different identities. Figure 5 also presents that when a small set of these distinctive features is removed (i.e., $x\cdot(1-m)$), the facial feature extractor struggles to discern these faces. That’s to say, if a feature extractor fails to extract the distinguished features of some facial images in the enrollment and authentication stage, these images will probably have higher similarities. That’s why many of us often encounter situations where we are mistakenly identified as other individuals.
Figure 5: Similarity distribution of the different mask regions. (MobileFace, CelebA)
Figure 6: Results of user study. Question 1:Which type of enrollment methods did you use? Question 2:Were you supervised during the enrollment process? Question 3:Were you supervised during the authentication process?
@@ 4 THREAT MODEL
However, managing this with only insider’s endeavor is not trivial. To achieve this, the attackers can provide the insider with a specific and wearable trigger used for disguise. The insider wearing a mask can enroll his/her face into the database. After that, both the insider and attackers can bypass the FRS by wearing the same trigger without the need for extra optimization during each authentication.
@@
Adversary’s Capabilities. In this paper, we assume that the compromised insider has the ability to enroll an identity in FRS. What needs to be emphasized is that this privilege is common for insiders of the FRS (e.g., students can update their facial images on a specified APP easily to avoid misidentification caused by their changes in appearance [13, 14]). They can complete this step by uploading or taking photos with IoT devices, and corresponding feature vectors will be stored in the database. It should be noted that both of these methods are practical in the real world. To demonstrate this, we expanded upon the user survey results from previous work [4] and added a question (73 users have answered Question 3 with details presented in Appendix B.4). Results shown in Figure 6 indicate that it is common for many FRS users to enroll either digital images or photos taken by edge devices and the devices used for evaluation also validate that both approaches are used in real-world scenarios (see in B.7). Moreover, FRS users are rarely supervised during both the enrollment and authentication phases. Thus, this paper considers two conditions: uploading facial images and taking photos. Uploading facial images directly means less digital-to-physical distortion while taking facial photos in the physical domain must consider this distortion, which will have a negative impact on FIBA.
@@ 5 METHODOLOGY
5.1 Attack Objective Design
Based on the preliminary analysis, attack objectives are proposed in the following paragraphs. Given a pretrained facial feature extractor $f_{\theta}$, an adversary intends to fool it by cooperating with a legitimate but compromised insider whose facial image $x_{v}$ can be enrolled in the feature database $\mathbb{D}$. Specifically, the adversary generates a wearable patch to replace the key features of the face and sends it to the insider. The insider wearing the patch then enrolls his/her face in the database. The adversary can also bypass the FRS by wearing the same patch. To this end, our optimization objective can be formalized as:
@@ 5.1 Attack Objective Design
where $p$ and $m$ denote patch and mask, respectively. Thus, $x_{v}\oplus p$ refers to the enrolled face with patch $p$ (backdoor trigger). $\mathcal{L}(\cdot)$ is the overall loss function.
@@ 5.2 Backdoor Trigger Enhancement
The key step in FIBA is to generate a backdoor trigger (known as a patch) $p$. An ideal generated patch should also have the following characteristics: 1) high generalization to other models since we cannot get access to FRS under the black-box setting; 2) high robustness to resist distortions made by many factors, e.g., illumination and lens resolution. To tackle this problem, we utilize several physical enhancement techniques presented in the following subsections.
Model Ensemble. Previous works have demonstrated the effectiveness of model ensemble [39] to enhance the generalization of the generated patch under the black-box setting. In this paper, we also adopt this technique while applying our attack to real-world scenarios. Specifically, we modified the $\mathcal{L}_{sim}$ as follows:
@@
Physical Transformation. As mentioned before, real-world loss should also be considered to mitigate the negative impact of cross-domain transformation. We construct a transformation set $\mathbb{T}$ to simulate this complex process to improve the patch’s robustness to digital-to-physical distortions. The first one is ”ColorJiggle”[44] which is used to adjust different brightness or hue changes when applied in real-world scenarios, corresponding to real-world lighting changes. ”RandomAffine” is applied to make the patch translation and rotation invariant, as we can not precisely adjust the disguised facial images to the corresponding position in the physical world. Additionally, motion blur, caused by relative motion between an object or scene and a camera during the exposure time of a photograph or the frame capture of a video, hinders the physical performance of FIBA. Here, we use ”RandomGaussianBlur” to reduce the influence of motion blur and printing errors introduced by the color printer. The pseudocode of the complete algorithm is provided in Appendix B.
@@ 6.1 Experimental Setup
Datasets. Digital experiments of feature extractors focus primarily on datasets: LFW [45], CALFW [46], AgeDB [47], and CelebA-HQ [35] (use CelebA for short). The details of the used datasets are presented in B.5.
All of these facial images are resized to 112×112 pixels to match the input requirements of models used for optimization. For experiments in the physical domain, facial images are taken from individuals using edge devices. For the training set of FIBA to generate the backdoor trigger in the real-world evaluation, we use 50 facial images (no overlap with images used for evaluation) from CelebA by default unless otherwise specified and the number of the training data is discussed in Appendix B.10.
@@
    Models with * denote white-box attack (source model is the same as target model).
Target Models, APIs and IoT Devices. We use six common facial feature extractors in our evaluation, with detailed information presented in Table X. Since the thresholds of these models are different, we unify the threshold ($\delta_{match}=0.35$) for simplicity, and a higher threshold also presents a more conservative attack performance and lower false positive rate (evaluation with higher $\delta_{match}$ can be found in Figure 7). To make the evaluation more practical, we also use five commercial APIs from different vendors: Aliyun (version: 20191230), Face++ (version: v4), ArcSoft (version: v3), Huawei (version: v2), and Tencent (version: v3). These vendors are selected since they are widely deployed and have become the mainstream of face recognition services. All of these APIs claim to achieve over 99% accuracy with less than 0.1% false positive rate. Edge devices are also considered since they are a common part of FRS. To simplify the expression, IoT devices from SenseTime, HIKVISION, and MoreDian are labeled as D-1, D-2, and D-3 for short, respectively. More details can be found in Table IX.
@@
$ASR=\frac{1}{N}\sum_{i=0}^{N}\mathbb{I}\left(\cos\left(f_{\theta}(x_{i}\oplus p),f_{\theta}(x_{v}\oplus p)\right)\geq\delta_{match}\right)$ (16)
@@ 6.2 Attack Performance
The results are presented in Table II. For comparison, the match rate of the benign faces is also given. Note that FIBA can achieve an ASR of 100% under a white-box setting even if the target dataset (first row for different source datasets) is not accessible. Under the black-box setting, where the victim model is not accessible, FIBA can still achieve remarkable performance with more than 99% for most combinations if CelebA is selected as the source dataset. When other datasets are chosen as the source dataset to train a backdoor trigger, there is a slight decrease in ASR. The observed decrease in ASR can be attributed to the fact that the other three datasets have a lower quality than CelebA, and only a small portion of these datasets is included in our training set. Consequently, the quality of the training set is pivotal to the optimization process of FIBA, and adversaries may feasibly gather such images (for instance, from CelebA) through artificial means. It’s also noteworthy that different models present different vulnerabilities under the black-box setting. CASIA-Softmax exhibits greater susceptibility, whereas MobilenetV2 demonstrates improved robustness against FIBA. Overall, FIBA proves to be effective and feasible for attackers, as no additional computation is needed once the backdoor trigger has been generated. The cross-model and cross-dataset results further demonstrate FIBA’s high generalization and universality in the digital domain.
@@
Despite minimal changes in the algorithm, a significant performance gap exists between FIBA and this adversarial attack. The disparity can be attributed to FIBA’s ability to exploit intrinsic flaws in FRS to substitute key facial features with generated trigger features. In contrast, adversarial patch-based attacks or previous enrollment-stage attacks [4, 12] have to balance the similarity trade-off of different identities. To make it clear, we visualize the facial features of the benign, triggered, and enrolled sample in Figure 15. Upon applying dimensionality reduction, it becomes evident that as the benign face image is masked with a backdoor trigger, its features intuitively converge with those of the enrolled sample. Additionally, we have discussed how different sizes of training data influence the performance of FIBA (details shown in Appendix B.10) and found out that even one image is enough to train a backdoor trigger with more than 90% ASR under black-box setting. This also distinguishes FIBA from universal adversarial attacks.
@@
Attack Performance with Different Threshold. In the previous evaluation, we set $\delta_{match}$ to 0.35. To comprehend this factor, we analyzed how $\delta_{match}$ affects the attack performance with the experimental results on CelebA (target dataset) shown in Figure 7. We can observe from the figure that most of the similarity under the black-box setting can achieve 0.7 since the ASR of most models can approach 100%. Consequently, FIBA can maintain a high ASR even when the threshold approaches 0.7, which also signifies the high transferability of FIBA across various models.
@@
API Vendors $\delta_{match}$ Similarity ASR (×$\times 100\%$)
@@ 6.3 Real-world Evaluation
API Evaluation in Digital Domain. We employ the model ensemble technique as mentioned in Section 5 to generate a backdoor trigger (50 images for training) and evaluate its efficacy with 500 samples from CelebA (different identities than those in the training set). The results are presented in Table V. Face similarity thresholds of different vendors’ FRS services and the average similarity of FIBA are also given in the table. In scenarios where the API returns ”no face detected” for a given sample, said sample is categorized as a failed sample. An attack is deemed successful solely if the response similarity exceeds the given threshold of the corresponding vendor. Note that for APIs from three vendors: ArcSoft, Alibaba, and Face++, FIBA is able to achieve nearly 100% ASR. Additionally, the similarity is much higher than $\delta_{match}$. However, for APIs from Huawei and Tencent, FIBA exhibits inferior attack performance. This could be attributed to these vendors having utilized certain detection methods or deployed robust models. Illustrations (not included in evaluation) of FIBA are depicted in Figure 9.
@@
The results are shown in Table VI. In contrast to digital authentication, both the ASR and the average similarity of FIBA have experienced declines to varying degrees, notably for the API from ”Alibaba”. However, FIBA can still attack ”ArcSoft” and ”Face++” with more than 90% ASR, and it’s unbearable for some safety-critical applications that have been equipped with these APIs. Also, there hasn’t been a big drop in ASR between the digital and physical domains, with even a slight increase in average similarity for several APIs. We suppose that’s because distortions between the physical and digital backdoor trigger counteracts the negative effect raised by physical authentication, which also validates the robustness of FIBA against physical distortion.
@@
Impact of Light Condition. Another important factor that needs to be studied in the physical domain is that the environment’s lighting conditions cannot be guaranteed and pose a significant challenge to existing attacks. To evaluate the impact of different lighting conditions on FIBA, we conducted experiments with varying lighting conditions, and all other settings are the same as those in the physical domain (physical enrollment and authentication with the ArcSoft API). More details are presented in Appendix LABEL:appendix:experiments.
@@
Impact of Shooting Angles. Apart from environmental factors such as lighting conditions, shooting angles also play a significant role during the authentication stage. Attacks that can only perform well with specific shooting angles are not applicable in the real world. To evaluate the performance of FIBA against different shooting angles in the physical domain, we conduct an experiment using ArcSoft API with the same settings as in the previous evaluation. More details are presented in Appendix B.11.

### v3

@@ 3.3 Experimental Exploration and Analysis
$\mathrm{cos}(f_{\theta}(\Phi(x_{q})),e_{q})\leq\mathrm{\delta_{match}}\quad or$ (3)
$\mathrm{cos}(f_{\theta}(\Phi(x_{q})),e_{k})\geq\mathrm{\delta_{match}},k\neq q$
where facial features $e_{q},e_{k}\in\mathbb{D}$ and $\Phi(\cdot)$ denotes the transformation that makes $f_{\theta}(\Phi(x_{q}))$ different from $e_{q}$, leading to the misidentification. In practice, $\Phi(\cdot)$ simulates the factors introduced by the environment (e.g., light intensity, quality of the images, etc) and human (makeup, shooting angles, expressions, decorations, etc). All of these factors work together and result in recognition errors.
@@
where $l(\cdot)$ is the loss function for optimization. We also restrict the changing area of $p$ with $\|p\|_{n}\leq\gamma$. The details of the algorithm can be found in Appendix 1. For the sake of experimental simplicity, we only set the changing areas of $p$ to 10% (for image space), and $p$ is randomly initialized in each iteration to avoid its influence on the optimization of $m$. The result is shown in Figure 4, using CelebA [34] and MobileFace [35] for optimization. From the similarity between $x$ and $x\oplus p$, we can indicate that some features are more essential for $f_{\theta}$ than others, and only a small part of the occlusion of these key features can blind FRS. Based on the analysis above, the unrecognition ($\mathrm{cos}(f_{\theta}(\Phi(x_{q})),e_{q})\leq\mathrm{\delta_{match}}$) of FRS can be attributed to the influence of various factors on the key features of feature extraction.
@@
The above inference can also be associated with the misidentification ($\mathrm{cos}(f_{\theta}(\Phi(x_{q})),e_{k})\geq\mathrm{\delta_{match}},k\neq q$) of FRS, since faces without distinguished features could easily be matched with an “average face” which also lacks distinguished features in the database. To validate this conjecture, we randomly sample 1000 face pairs $(x_{i},x_{j})$ from CelebA as $\mathbb{D}_{pair}$, where the identity of $x_{i}$ is different from that of $x_{j}$. Then, we use prefabricated masks (see in Appendix 1) $\mathbb{M}=\{m_{0},m_{1},...,m_{i}\}$ for different parts of the face. Based on the above preparations, we can draw a histogram of the similarity of $\mathrm{cos}(f_{\theta}(x_{i}\cdot(1-m_{k})),f_{\theta}(x_{j}\cdot(1-m_{k})))$, for $m_{k}\in\mathbb{M}$ and $(x_{i},x_{j})\in\mathbb{D}_{pair}$. Figure 5 illustrates that the distribution of face similarity will move forward as a result of the addition of the masks, and different regions present different impacts on the similarity changes. For instance, mask regions, including the eye and nose, can largely increase the similarity of faces from different identities. However, masks of mouth or cheek have less impact on the feature extraction of FRS. With these observations, we can also arrive at a conclusion consistent with the previous experiments, which is that facial feature extractors tend to focus more on the local features of faces to distinguish between facial images from different identities. Figure 5 also presents that when a small set of these distinctive features is removed (i.e., $x\cdot(1-m)$), the facial feature extractor struggles to discern these faces. That’s to say, if a feature extractor fails to extract the distinguished features of some facial images in the enrollment and authentication stage, these images will probably have higher similarities. That’s why many of us often encounter situations where we are mistakenly identified as other individuals.
Figure 5: Similarity distribution of the different mask regions (on CelebA dataset with MobileFace).
Figure 6: Results of user study. From left to right are, Question 1:Which type of enrollment did you use often? Question 2:Were you supervised during the enrollment? Question 3:Were you supervised during the authentication?
@@ 4 THREAT MODEL
However, managing this with only an insider’s endeavor is not trivial. To achieve this, the attackers can provide the insider with a specific and wearable trigger used for disguise. The insider wearing a mask can enroll his/her face into the database. After that, both the insider and attackers can bypass the FRS by wearing the same trigger without the need for extra optimization during each authentication.
@@
Adversary’s Capabilities. In this paper, we assume that the compromised insider has the ability to enroll an identity in FRS. What needs to be emphasized is that this privilege is common for insiders of the FRS (e.g., students can update their facial images on a specified APP easily to avoid misidentification caused by their changes in appearance [13, 12]). They can complete this step by uploading or taking photos with IoT devices, and corresponding feature vectors will be stored in the database. It should be noted that both of these methods are practical in the real world. To demonstrate this, we expanded upon the user survey results from a previous work [4] and added a question (73 users have answered Question 3 with details presented in Appendix 5). Results shown in Figure 6 indicate that it is common for many FRS users to enroll either digital images or photos taken by edge devices and the devices used for evaluation also validate that both approaches are used in real-world scenarios (see in 8). Moreover, FRS users are rarely supervised during both the enrollment and authentication phases. Thus, this paper considers two conditions: uploading facial images and taking photos. Uploading facial images directly means less digital-to-physical distortion while taking facial photos in the physical domain must consider this distortion, which will have a negative impact on FIBA.
@@ 5 METHODOLOGY
5.1 Backdoor Trigger Generation
Based on the preliminary analysis, attack objectives are proposed in the following paragraphs. Given a pretrained facial feature extractor $f_{\theta}$, an adversary intends to fool it by cooperating with a legitimate but compromised insider whose facial image $x_{v}$ can be enrolled in the feature database $\mathbb{D}$. Specifically, the adversary generates a wearable patch to replace the key features of the face and sends it to the insider. The insider wearing the patch then enrolls his/her face in the database. The adversary wearing the same patch can also bypass the FRS. To this end, our optimization objective can be formalized as:
@@ 5.1 Backdoor Trigger Generation
where $p$ and $m$ denote patch and mask, respectively. Thus, $x_{v}\oplus p$ refers to the enrolled face with patch $p$ (backdoor trigger). $\mathcal{L}(\cdot)$ is the overall loss to achieve adversary’s goal.
@@ 5.2 Backdoor Trigger Enhancement
The key step in FIBA is to generate a backdoor trigger (known as a patch) $p$. An ideal generated patch should also have the following characteristics: 1) high generalization to other models since we cannot get access to FRS under the black-box setting; 2) high robustness to resist distortions made by many factors, e.g., illumination and lens resolution. To achieve this, we utilize several physical enhancement techniques presented in the following subsections.
Model Ensemble. Previous works have demonstrated the effectiveness of model ensemble [38] in augmenting the generalization of the generated patch within the black-box setting. In this paper, we also adopt this technique while applying our attack to real-world scenarios. Specifically, we modify the $\mathcal{L}_{sim}$ as follows:
@@
Physical Transformation. As mentioned before, real-world loss should also be considered to mitigate the negative impact of cross-domain transformation. We construct a transformation set $\mathbb{T}$ to simulate this complex process to improve the patch’s robustness to digital-to-physical distortions. The first one is “ColorJiggle”[43] which is used to adjust different brightness or hue changes when applied in real-world scenarios, corresponding to real-world lighting changes. “RandomAffine” is applied to make the patch translation and rotation invariant, as we can not precisely adjust the disguised facial images to the corresponding position in the physical world. Additionally, motion blur, caused by relative motion between an object or scene and a camera during the exposure time of a photograph or the frame capture of a video, hinders the physical performance of FIBA. Here, we use “RandomGaussianBlur” to reduce the influence of motion blur and printing errors introduced by the color printer. The pseudocode of the complete algorithm is provided in Appendix 2.
@@ 6.1 Experimental Setup
Datasets. Digital experiments of feature extractors focus primarily on datasets: LFW [44], CALFW [45], AgeDB [46], and CelebA-HQ [34] (use CelebA for short). The details of the used datasets are presented in Appendix 6.
All of these facial images are resized to 112×112 pixels to match the input requirements of models used for optimization. For experiments in the physical domain, facial images are taken from individuals using edge devices. For the training set of FIBA to generate the backdoor trigger in the real-world evaluation, we use 50 facial images (no overlap with images used for evaluation) from CelebA by default unless otherwise specified and the number of the training data is discussed in Appendix 11.
@@
    Models with * denote the white-box attack (source model is the same as the target model).
Target Models, APIs and IoT Devices. We use six common facial feature extractors in our evaluation, with detailed information presented in Table 3. Since the thresholds of these models are different, we unify the threshold ($\mathrm{\delta_{match}}=0.35$) for simplicity, and a higher threshold also presents a more conservative attack performance and lower false positive rate (evaluation with higher $\mathrm{\delta_{match}}$ can be found in Figure 7). To make the evaluation more practical, we also use five commercial APIs from different vendors: Aliyun (version: 20191230), Face++ (version: v4), ArcSoft (version: v3), Huawei (version: v2), and Tencent (version: v3). These vendors are selected since they are widely deployed and have become the mainstream of face recognition services. All of these APIs claim to achieve over 99% accuracy with less than 0.1% false positive rate. Edge devices are also considered since they are a common part of FRS. To simplify the expression, IoT devices from SenseTime, HIKVISION, and MoreDian are labeled as D-1, D-2, and D-3 for short, respectively. More details can be found in Table 2.
@@
$ASR=\frac{1}{N}\sum_{i=0}^{N}\mathbb{I}\left(\cos\left(f_{\theta}(x_{i}\oplus p),f_{\theta}(x_{v}\oplus p)\right)\geq\mathrm{\delta_{match}}\right)$ (16)
@@ 6.2 Attack Performance
The results are presented in Table II. For comparison, the match rate of the benign faces is also given. Note that FIBA can achieve an ASR of 100% under a white-box setting even if the target dataset (first row for different source datasets) is inaccessible. Under the black-box setting, where the victim model is inaccessible, FIBA can still achieve remarkable performance with more than 99% for most combinations if CelebA is selected as the source dataset. When other datasets are chosen as the source dataset to train a backdoor trigger, there is a slight decrease in ASR. The observed decrease in ASR can be attributed to the fact that the other three datasets have a lower quality than CelebA, and only a small portion of these datasets is included in our training set. Consequently, the quality of the training set is pivotal to the optimization process of FIBA, and adversaries may feasibly gather such images (for instance, from CelebA) through artificial means. It’s also noteworthy that different models present different vulnerabilities under the black-box setting. CASIA-Softmax exhibits greater susceptibility, whereas MobilenetV2 demonstrates improved robustness against FIBA. Overall, FIBA proves to be effective and feasible for attackers, as no additional computation is needed once the backdoor trigger has been generated. The cross-model and cross-dataset results further demonstrate FIBA’s high generalization and universality in the digital domain.
@@
As shown in Table III, there is a significant performance gap exists between FIBA and this adversarial attack. The disparity can be attributed to FIBA’s ability to exploit intrinsic flaws in FRS to substitute key facial features with generated trigger features. In contrast, adversarial patch-based attacks or previous enrollment-stage attacks [4, 14] have to balance the similarity trade-off of different identities. To make it clear, we visualize the facial features of the benign, triggered, and enrolled sample in Figure 3. Upon applying dimensionality reduction, it becomes evident that as the benign face image is masked with a backdoor trigger, its features intuitively converge with those of the enrolled sample. Additionally, we have discussed how different sizes of training data influence the performance of FIBA (details shown in Appendix 11) and found out that even one image is enough to train a backdoor trigger with more than 90% ASR within a black-box setting. This also distinguishes FIBA from universal adversarial attacks.
@@
Attack Performance with Different Threshold. In the previous evaluation, we set $\mathrm{\delta_{match}}$ to 0.35. To comprehend this factor, we analyzed how $\mathrm{\delta_{match}}$ affects the attack performance with the experimental results on CelebA (target dataset) shown in Figure 7. We can observe from the figure that most of the similarity under the black-box setting can achieve 0.7 since the ASR of most models can approach 100%. Consequently, FIBA can maintain a high ASR even when the threshold approaches 0.7, which also signifies the high transferability of FIBA across various models.
@@
API Vendors $\mathrm{\delta_{match}}$ Similarity ASR (×$\times 100\%$)
@@ 6.3 Real-world Evaluation
API Evaluation in Digital Domain. We employ the model ensemble technique as mentioned in Section 5 to generate a backdoor trigger (50 images for training) and evaluate its efficacy with 500 samples from CelebA (different identities than those in the training set). The results are presented in Table V. Face similarity thresholds of different vendors’ FRS services and the average similarity of FIBA are also given in the table. In scenarios where the API returns “no face detected” for a given sample, said sample is categorized as a failed sample. An attack is deemed successful solely if the response similarity exceeds the given threshold of the corresponding vendor. Note that for APIs from three vendors: ArcSoft, Alibaba, and Face++, FIBA is able to achieve nearly 100% ASR. Additionally, the similarity is much higher than $\mathrm{\delta_{match}}$. However, for APIs from Huawei and Tencent, FIBA exhibits inferior attack performance. This could be attributed to these vendors having utilized certain detection methods or deployed robust models. Illustrations (not included in evaluation) of FIBA are depicted in Figure 9.
@@
The results are shown in Table VI. In contrast to digital authentication, both the ASR and the average similarity of FIBA have experienced declines to varying degrees, notably for the API from “Alibaba”. However, FIBA can still attack “ArcSoft” and “Face++” with more than 90% ASR, and it’s unbearable for some safety-critical applications that have been equipped with these APIs. Also, there hasn’t been a big drop in ASR between the digital and physical domains, with even a slight increase in average similarity for several APIs. We suppose that’s because distortions between the physical and digital backdoor trigger counteracts the negative effect raised by physical authentication, which also validates the robustness of FIBA against physical distortion.
@@
Impact of Light Condition. Another important factor that needs to be studied in the physical domain is that the environment’s lighting conditions cannot be guaranteed and pose a significant challenge to existing attacks. To evaluate the impact of different lighting conditions on FIBA, we conducted experiments with varying lighting conditions, and all other settings are the same as those in the physical domain (physical enrollment and authentication with the ArcSoft API). More details are presented in Appendix 12.
@@
Impact of Shooting Angles. Apart from environmental factors such as lighting conditions, shooting angles also play a significant role during the authentication stage. Attacks that can only perform well with specific shooting angles are not applicable in the real world. To evaluate the performance of FIBA against different shooting angles in the physical domain, we conduct an experiment using ArcSoft API with the same settings as in the previous evaluation. More details are presented in Appendix 12.